OpenStudy (anonymous):

Simplify x^2+8x/x^2 A x+8/x^2 B x+8/x C 8x D 8

OpenStudy (anonymous):

C?

OpenStudy (saifoo.khan):

how

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

idk

OpenStudy (safari321):

yeah explain

OpenStudy (anonymous):

it will be 1 + 8/x

OpenStudy (safari321):

why you think that

OpenStudy (saifoo.khan):

\[\frac{x(x+8)}{x \times x}=\frac{\cancel x(x+8)}{\cancel x \times x}=\frac{x+8}{x}\]

OpenStudy (anonymous):

x/x=1